# Presentations of Schur and Specht modules in characteristic zero

### v1

@@
Abstract.
New presentations of Specht modules of symmetric groups over fields of characteristic zero have been obtained by Brauner, Friedmann, Hanlon, Stanley and Wachs. These involve generators that are column tabloids and relations that are symmetrizations of Garnir relations with minimal number of exchanges between consecutive columns or Garnir relations with maximal number of exchanges between consecutive columns. In this paper, we examine symmetrizations of Garnir relations and Garnir relations with any number of exchanges. In both cases, we provide sufficient arithmetic conditions so that the corresponding quotient is a Specht module. In particular, in the second case this yields new presentations of Specht modules if the parts of the conjugate partition that correspond to maximal number of exchanges greater than 1 are distinct. These results generalize the presentations mentioned above and offer an answer to a question of Friedmann, Hanlon and Wachs. Our approach is via representations of the general linear group.
Key words and phrases:
Specht module, Schur module, column tabloid, Garnir relations
1991 Mathematics Subject Classification:
Primary 05E10, 20C30; Secondary 20G05
1. Introduction
@@
(1.1) $g_{c,k}^{T}=\overline{T}-\sum\overline{S},$
where the sum is over all $S\in\mathcal{T}_{\lambda}$ obtained from $T$ by exchanging the top $k$ entries of column $c+1$ with any $k$ entries from column $c$ of $T$ while maintaining the vertical order of each of the exchanged sets. Let $G^{\lambda}$ be the subspace of $M^{\lambda}$ spanned by all $g_{c,k}^{T}$ as $T,c,k$ vary. (In [4], the notation $Q^{\lambda}$ is used.) Then $G^{\lambda}$ is an $\mathfrak{S}_{r}$-submodule of $M^{\lambda}$ and we have the important fact that $M^{\lambda}/G^{\lambda}\simeq S^{\lambda}$, see [4, p. 99]. It is well known that, since the characteristic of $K$ is equal to zero, $G^{\lambda}$ is spanned by the $g_{c,1}^{T}$ as $T,c$ vary.
1.1. Motivation
@@
(1.2) $h_{c}^{T}=\lambda_{c+1}^{\prime}\overline{T}-\sum\overline{S},$
where the sum ranges over all tableaux $S$ obtained from $T$ by exchanging one entry in column $c+1$ of $T$ with one entry in column $c$. Let $G^{\lambda,min}$ be the subspace of $M^{\lambda}$ spanned by all $h_{c}^{T}$, where $T\in\mathcal{T}_{\lambda}$ and $c$ is a column of $T$. (In [1], $G^{\lambda,min}$ is denoted by $\tilde{H}_{\lambda}$.) The following result was shown in [1].
@@
If $\lambda$ is a partition of $r$, then as $\mathfrak{S}_{r}$-modules we have $M^{\lambda}/G^{\lambda,min}\simeq S^{\lambda}.$
1.2. Results
@@
(1.3) $r^{T}_{c,k_{c}}=\tbinom{\lambda_{c+1}^{\prime}}{k_{c}}\overline{T}-\sum\overline{S},$
where the sum is over all $S\in\mathcal{T}_{\lambda}$ obtained from $T$ by exchanging any $k_{c}$ entries of column $c+1$ with any $k_{c}$ entries from column $c$ of $T$ while maintaining the vertical order of each of the exchanged sets. Let $SGR^{\lambda}(k_{1},\dots,k_{{\lambda_{1}}-1})$ be the subspace of $M^{\lambda}$ spanned by all $r^{T}_{c,k_{c}}$, where $T\in\mathcal{T}_{\lambda}$ and $c$ is a column of $T$ (symmetrized Garnir relations). In particular, we have two extreme cases; if $k_{1}=\dots=k_{{\lambda_{1}}-1}=1$, then $SGR^{\lambda}(1,\dots,1)=G^{\lambda,min}$, and if $k_{1}=\lambda^{\prime}_{2},\dots,k_{{\lambda_{1}}-1}=\lambda^{\prime}_{\lambda_{1}}$, then $SGR^{\lambda}(\lambda^{\prime}_{2},\dots,\lambda^{\prime}_{\lambda_{1}})=G^{\lambda,max}$. The subspace $SGR^{\lambda}(k_{1},\dots,k_{\lambda_{1}-1})$ is an $\mathfrak{S}_{r}$-submodule of $M^{\lambda}$. In Theorem 6.2, we obtain sufficient arithmetic conditions on $\lambda$ and $k_{1},\dots,k_{{\lambda_{1}}-1}$ so that $M^{\lambda}/SGR^{\lambda}(k_{1},\dots,k_{{\lambda_{1}}-1})\simeq S^{\lambda}$.
@@
Let $\lambda$ be a partition of $r$. Then as $\mathfrak{S}_{r}$-modules we have
$M^{\lambda}/SGR^{\lambda}(k_{1},\dots,k_{\lambda_{1}-1})\simeq S^{\lambda}$ if
@@
for all $c=1,\dots,\lambda_{1}-1$ and $j=1,\dots,k_{c}$.
@@
2. Preliminaries
@@
The Schur module of $G$ corresponding to a partition $\mu$ will be denoted $L_{\mu}$, see [9, 2.1]. This is defined as a quotient of $\Lambda^{\mu}$. Let us denote the corresponding projection by $\pi_{\mu}:\Lambda^{\mu}\to L_{\mu}$. With respect to the action of the diagonal matrices in $G$, $L_{\mu}$ has highest weight $\mu^{\prime}$, the conjugate partition of $\mu$. The formal character $ch(L_{\mu})$ of $L_{\mu}$ is the classical Schur polynomial $s_{\mu^{\prime}}$, [4, 2.2]. For example, if $\mu=(a)$ consists of one part, then $ch(L_{(a)})$ is $s_{(1^{a})}$, the elementary symmetric polynomial of degree $a$. We note that our $L_{\mu}$ is isomorphic to $E^{\lambda}$ of [4], where $\lambda=\mu^{\prime}$.
@@
(2.1) $\sum_{u,v}\sum_{t_{1}+t_{2}=t}(-1)^{(a_{2}-t_{2})t_{1}}x_{u}(a_{1}-t_{1})y_{v}(a_{2}-t_{2})\otimes x^{\prime}_{u}(t_{1})y^{\prime}_{v}(t_{2}),$
@@
{tikzcd}{tikzcd}
@@
(2.2) $\delta_{s}\circ\delta_{t}=\tbinom{s+t}{t}\delta_{s+t}\;\text{and}\;\theta_{s}\circ\theta_{t}=\tbinom{s+t}{t}\theta_{s+t}.$
@@
3. Presentations of Schur modules
@@
3.1. Two lemmas
We will need the relations of $L_{(a,b)}$ described in the following lemma, which is an analog of [7, Lemma 4.2] for Schur modules.
@@ 3.1. Two lemmas
(3.1) $(-1)^{a_{2}+a_{3}+1}\frac{1}{b_{1}}\big{(}(-1)^{a_{3}+b_{1}-1}(b_{2}+1)\pi_{\alpha}(A)+(-1)^{b_{1}+b_{2}-1}(b_{3}+1)\pi_{\alpha}(B)\big{)},$
@@
Let $a,b,i,t\in\mathbb{N}$ such that $a\geq b\geq i$ and let $\mu=(a+i,b-i)$. (1) Let $i<t$. Then the restriction of $\Phi_{t}$ to the summand $L_{\mu}$ of $\Lambda^{a}\otimes\Lambda^{b}$ is equal to zero. (2) Let $i\geq t$. Define $c_{t}(a,b,i)=\tbinom{i}{t}\tbinom{a-b+t+i}{t}$. Then the following diagram is commutative
{tikzcd}
@@
(1) By definition, $\Phi_{t}$ factors through $\Lambda^{a+t}\otimes\Lambda^{b-t}$. But the irreducible module $L_{\mu}$ is not a summand of $\Lambda^{a+t}\otimes\Lambda^{b-t}$ if $i<t$, according to Pieri’s rule. (2) For the first statement, it suffices to show that in the following diagram the left and right rectangles commute
{tikzcd}
The left rectangle commutes by the first equality of (2.2). Let $x\otimes y\in\Lambda^{a+t}\otimes\Lambda^{b-t}$. In the clockwise direction we have
@@
By the first identity of [7, Lemma 4.1(2)], we have $\sum_{i_{1}=0}^{t}(-1)^{t-i_{1}}\tbinom{a+i_{1}}{i_{1}}\tbinom{b-i}{t-i_{1}}=$ $\tbinom{a-b+t+i}{t}$. Hence $\pi_{\mu}\circ\delta_{i}\circ\theta_{t}(x\otimes y)$ is equal to
@@
and the right rectangle commutes. Since $n\geq a+b$, we have the element $x=e_{1}\cdots e_{a}\otimes e_{a+1}\cdots e_{a+b}\in\Lambda^{a}\otimes\Lambda^{b}$. From the definition of $\delta_{i}$ it follows that $\pi_{\mu}\circ\delta_{i}(x)$ is a sum of elements of the form ±$\pm\pi_{\mu}(e_{1}\cdots e_{a}u_{1}\cdots u_{i}\otimes u_{i+1}\cdots u_{b}),$ where $u_{1},\dots,u_{b}\in\{e_{a+1},\dots,e_{b}\}$ and $\{u_{1},\dots,u_{i}\}\cap$$\{u_{i+1},\dots,u_{b}\}=\emptyset$. The number of the $u_{i+1},\dots,u_{b}$ is equal to $b-i$ and we have $b-i\leq a.$ Thus there are no violations in the columns, meaning that the above elements are (up to sign) cosemistandard basis elements of $L_{\mu}$. They are clearly distinct. Hence we conclude that $\pi_{\mu}\circ\delta_{i}(x)$ is nonzero and the map $\pi_{\mu}\circ\delta_{i}$ is nonzero. ∎
@@ 3.2. Main result
From Remark 3.4 it follows that $\operatorname{Im}(\sum_{c=1}^{{\ell}-1}\overline{\Psi}_{\mu(c)})\subseteq\operatorname{Im}\theta_{\mu}$, where $\theta_{\mu}$ is the map in Theorem 2.2. Hence $\operatorname{Coker}(\sum_{c=1}^{{\ell}-1}\overline{\Psi}_{\mu(c)})\simeq L_{\mu}$ if and only if $\operatorname{Im}(\sum_{c=1}^{{\ell}-1}\overline{\Psi}_{\mu(c)})=\operatorname{Im}\theta_{\mu}$. Again by Remark 3.4,
(3.2) $\operatorname{Im}({\Psi}_{\mu(c)})\subseteq\operatorname{Im}(\Lambda^{\mu_{c}+1}\otimes\Lambda^{\mu_{c+1}-1}\xrightarrow{\theta_{1}}\Lambda^{\mu_{c}}\otimes\Lambda^{\mu_{c+1}}).$
@@
4. Schur modules and symmetrization of Garnir relations
@@
It is straightforward to verify that if $x\otimes y=e_{i_{1}}\cdots e_{i_{a}}\otimes e_{j_{1}}\cdots e_{j_{b}}$ is a basis element of $\Lambda^{a}\otimes\Lambda^{b}$, then $\phi_{k}(x\otimes y)$ is obtained by exchanging any $k$ elements $x_{1}<x_{2}<\dots<x_{k}$ of $\{e_{i_{1}},\dots,e_{i_{a}}\}$ with any $k$ elements $y_{1}<y_{2}<\dots<y_{k}$ of $\{e_{j_{1}},\dots,e_{j_{b}}\}$ and putting the $x_{1},\dots,x_{k}$ (respectively, $y_{1},\dots,y_{k}$) in the original positions of the $y_{1},\dots,y_{k}$ (respectively, $x_{1},\dots,x_{k}$). where
@@
(4.1) $\Phi_{t}=\sum_{i=0}^{t}(-1)^{t-i}\tbinom{b-t+i}{i}\phi_{t-i}=\sum_{k=0}^{t}(-1)^{k}\tbinom{b-k}{b-t}\phi_{k},$
where in the last equality we set $k=t-i$. Consider the $(b+1)\times(b+1)$ matrix $A=(a_{kt})$, where $a_{kt}=(-1)^{k}\tbinom{b-k}{b-t}$ and $0\leq k,t\leq b$. From the binomial identity $\sum_{j=0}^{b}(-1)^{k+j}\tbinom{b-k}{b-j}\tbinom{b-j}{b-t}=\delta_{k,t}$, see [8, p. 4], where $\delta_{k,t}$ is Kronecker’s delta, it follows that $A^{2}=I$, the identity matrix. Hence we may invert (4.1) to obtain $\phi_{k}=\sum_{t=0}^{k}(-1)^{t}\tbinom{b-t}{b-k}\Phi_{t}$ as desired. The second equality follows from the first and Definition 4.1. ∎
@@
for all $j=1,\dots,k$.
5. Schur modules and Garnir relations

### v3

@@
Abstract
New presentations of Specht modules of symmetric groups over fields of characteristic zero have been obtained by Brauner, Friedmann, Hanlon, Stanley and Wachs. These involve generators that are column tabloids and relations that are Garnir relations with maximal number of exchanges between consecutive columns or symmetrization of Garnir relations with minimal number of exchanges between consecutive columns. In this paper, we examine Garnir relations and their symmetrization with any number of exchanges. In both cases, we provide sufficient arithmetic conditions so that the corresponding quotient is a Specht module. In particular, in the first case this yields new presentations of Specht modules if the parts of the conjugate partition that correspond to maximal number of exchanges greater than 1 are distinct. These results generalize the presentations mentioned above and offer an answer to a question of Friedmann, Hanlon and Wachs. Our approach is via representations of the general linear group.
keywords:
Schur module , Specht module , Garnir relations , symmetric group
MSC:
[2020]20C30, 05E10, 20G05
1 Introduction
@@
$g_{c,k}^{T}=\overline{T}-\sum\overline{S},$ (1.1)
where the sum is over all $S\in\mathcal{T}_{\lambda}$ obtained from $T$ by exchanging the top $k$ entries of column $c+1$ with any $k$ entries from column $c$ of $T$ while maintaining the vertical order of each of the exchanged sets. Let $G^{\lambda}$ be the subspace of $M^{\lambda}$ spanned by all $g_{c,k}^{T}$ as $T,c,k$ vary. (In [4], the notation $Q^{\lambda}$ is used.) Then $G^{\lambda}$ is an $\mathfrak{S}_{r}$-submodule of $M^{\lambda}$ and we have the important fact that $M^{\lambda}/G^{\lambda}\simeq S^{\lambda}$, see [4, p. 99]. Since the characteristic of $K$ is equal to zero, $G^{\lambda}$ is spanned by the $g_{c,1}^{T}$ as $T,c$ vary [4, p. 102].
1.1 Motivation
@@
$h_{c}^{T}=\lambda_{c+1}^{\prime}\overline{T}-\sum\overline{S},$ (1.2)
where the sum ranges over all tableaux $S$ obtained from $T$ by exchanging one entry in column $c+1$ of $T$ with one entry in column $c$. Let $SGR^{\lambda,min}$ be the subspace of $M^{\lambda}$ spanned by all $h_{c}^{T}$, where $T\in\mathcal{T}_{\lambda}$ and $c$ is a column of $T$. (In [1], $SGR^{\lambda,min}$ is denoted by $\tilde{H}_{\lambda}$.) The following result was shown in [1].
@@
If $\lambda$ is a partition of $r$, then as $\mathfrak{S}_{r}$-modules we have $M^{\lambda}/SGR^{\lambda,min}\simeq S^{\lambda}.$
1.2 Results
@@
$r^{T}_{c,k_{c}}=\tbinom{\lambda_{c+1}^{\prime}}{k_{c}}\overline{T}-\sum\overline{S},$ (1.3)
where the sum is over all $S\in\mathcal{T}_{\lambda}$ obtained from $T$ by exchanging any $k_{c}$ entries of column $c+1$ with any $k_{c}$ entries from column $c$ of $T$ while maintaining the vertical order of each of the exchanged sets. Let $SGR^{\lambda}(k_{1},\dots,k_{{\lambda_{1}}-1})$ be the subspace of $M^{\lambda}$ spanned by all $r^{T}_{c,k_{c}}$, where $T\in\mathcal{T}_{\lambda}$ and $c$ is a column of $T$ (symmetrized Garnir relations). In particular, we have two extreme cases; if $k_{1}=\dots=k_{{\lambda_{1}}-1}=1$, then $SGR^{\lambda}(1,\dots,1)=SGR^{\lambda,min}$, and if $k_{1}=\lambda^{\prime}_{2},\dots,k_{{\lambda_{1}}-1}=\lambda^{\prime}_{\lambda_{1}}$, then $SGR^{\lambda}(\lambda^{\prime}_{2},\dots,\lambda^{\prime}_{\lambda_{1}})=G^{\lambda,max}$. The subspace $SGR^{\lambda}(k_{1},\dots,k_{\lambda_{1}-1})$ is an $\mathfrak{S}_{r}$-submodule of $M^{\lambda}$. In Theorem 6.23, we obtain sufficient arithmetic conditions on $\lambda$ and $k_{1},\dots,k_{{\lambda_{1}}-1}$ so that $M^{\lambda}/SGR^{\lambda}(k_{1},\dots,k_{{\lambda_{1}}-1})\simeq S^{\lambda}$.
@@
Let $\lambda$ be a partition of $r$ and let $\mu$ be the conjugate partition of $\lambda$. Then as $\mathfrak{S}_{r}$-modules we have $M^{\lambda}/SGR^{\lambda}(k_{1},\dots,k_{\lambda_{1}-1})\simeq S^{\lambda}$ if
@@
for all $c=1,\dots,\lambda_{1}-1$ and $j=1,\dots,\mu_{c+1}$.
@@
Remark.
This version of the paper differs from the published version as follows. A typo in the statement of Corollary 4.4 and Corollary 6.1 has been corrected, instead of $j=1,\dots,b$, it should be $j=1,\dots,k$. Also, in the statement of Theorem 6.2 it should be $j=1,\dots,\mu_{c+1}$. Examples 6.3(1) and 6.3(3) have been corrected accordingly. We are indebted to Tamar Friedmann for pointing out these corrections.
2 Preliminaries
@@
The Schur module of $G$ corresponding to a partition $\mu$ will be denoted $L_{\mu}$, see [10, 2.1]. This is defined as a quotient of $\Lambda^{\mu}$. Let us denote the corresponding projection by $\pi_{\mu}:\Lambda^{\mu}\to L_{\mu}$. With respect to the action of the diagonal matrices in $G$, $L_{\mu}$ has highest weight $\mu^{\prime}$, the conjugate partition of $\mu$. The formal character $ch(L_{\mu})$ of $L_{\mu}$ is the classical Schur polynomial $s_{\mu^{\prime}}$, [4, 2.2]. For example, if $\mu=(a)$ consists of one part, then $ch(L_{(a)})$ is $s_{(1^{a})}$, the elementary symmetric polynomial of degree $a$. We note that our $L_{\mu}$ is isomorphic to $E^{\lambda}$ of [4], where $\lambda=\mu^{\prime}$.
@@
$\sum_{u,v}\sum_{t_{1}+t_{2}=t}(-1)^{(a_{2}-t_{2})t_{1}}x_{u}(a_{1}-t_{1})y_{v}(a_{2}-t_{2})\otimes x^{\prime}_{u}(t_{1})y^{\prime}_{v}(t_{2}),$ (2.1)
@@
${\Lambda^{a+t}\otimes\Lambda^{b-t}}$${\Lambda^{a}\otimes\Lambda^{b}}$${\Lambda^{a+t+s}\otimes\Lambda^{b-t-s}}$$\scriptstyle{\delta_{s}}$$\scriptstyle{\delta_{t}}$$\scriptstyle{\delta_{t+s}}$${\Lambda^{a-t}\otimes\Lambda^{b+t}}$${\Lambda^{a}\otimes\Lambda^{b}}$${\Lambda^{a-t-s}\otimes\Lambda^{b+t+s}}$$\scriptstyle{\theta_{s}}$$\scriptstyle{\theta_{t}}$$\scriptstyle{\theta_{t+s}}$
@@
$\delta_{s}\circ\delta_{t}=\tbinom{s+t}{t}\delta_{s+t}\;\text{and}\;\theta_{s}\circ\theta_{t}=\tbinom{s+t}{t}\theta_{s+t}.$ (2.2)
@@
3 Presentations of Schur modules
@@
3.1 Two lemmas
We will need the relations of $L_{(a,b)}$ described in the following lemma, which is an analog of [8, Lemma 4.2] for Schur modules.
@@ 3.1 Two lemmas
$(-1)^{a_{2}+a_{3}+1}\frac{1}{b_{1}}\big{(}(-1)^{a_{3}+b_{1}-1}(b_{2}+1)\pi_{\alpha}(A)+(-1)^{b_{1}+b_{2}-1}(b_{3}+1)\pi_{\alpha}(B)\big{)},$ (3.1)
@@
Let $a,b,i,t\in\mathbb{N}$ such that $a\geq b\geq i$ and let $\mu=(a+i,b-i)$.
(1) Let $i<t$. Then the restriction of $\Phi_{t}$ to the summand $L_{\mu}$ of $\Lambda^{a}\otimes\Lambda^{b}$ is equal to zero.
(2) Let $i\geq t$. Define $c_{t}(a,b,i)=\tbinom{i}{t}\tbinom{a-b+t+i}{t}$. Then the following diagram is commutative
[…]
@@
(1) By definition, $\Phi_{t}$ factors through $\Lambda^{a+t}\otimes\Lambda^{b-t}$. But the irreducible module $L_{\mu}$ is not a summand of $\Lambda^{a+t}\otimes\Lambda^{b-t}$ if $i<t$, according to Pieri’s rule.
(2) For the first statement, it suffices to show that in the following diagram the left and right rectangles commute
[…]
The left rectangle commutes by the first equality of (2.2).
Let $x\otimes y\in\Lambda^{a+t}\otimes\Lambda^{b-t}$. In the clockwise direction we have
@@
By the first identity of [8, Lemma 4.1(2)], we have $\sum_{i_{1}=0}^{t}(-1)^{t-i_{1}}\tbinom{a+i_{1}}{i_{1}}\tbinom{b-i}{t-i_{1}}=$ $\tbinom{a-b+t+i}{t}$. Hence $\pi_{\mu}\circ\delta_{i}\circ\theta_{t}(x\otimes y)$ is equal to
@@
and the right rectangle commutes.
Since $n\geq a+b$, we have the element $x=e_{1}\cdots e_{a}\otimes e_{a+1}\cdots e_{a+b}\in\Lambda^{a}\otimes\Lambda^{b}$. From the definition of $\delta_{i}$ it follows that $\pi_{\mu}\circ\delta_{i}(x)$ is a sum of elements of the form ±$\pm\pi_{\mu}(e_{1}\cdots e_{a}u_{1}\cdots u_{i}\otimes u_{i+1}\cdots u_{b}),$ where $u_{1},\dots,u_{b}\in\{e_{a+1},\dots,e_{b}\}$ and $\{u_{1},\dots,u_{i}\}\cap$$\{u_{i+1},\dots,u_{b}\}=\emptyset$. The number of the $u_{i+1},\dots,u_{b}$ is equal to $b-i$ and we have $b-i\leq a.$ Thus there are no violations in the columns, meaning that the above elements are (up to sign) cosemistandard basis elements of $L_{\mu}$. They are clearly distinct. Hence we conclude that $\pi_{\mu}\circ\delta_{i}(x)$ is nonzero and the map $\pi_{\mu}\circ\delta_{i}$ is nonzero. ∎
@@ 3.2 Main result
From Remark 3.10 it follows that $\operatorname{Im}(\sum_{c=1}^{{\ell}-1}\overline{\Psi}_{\mu(c)})\subseteq\operatorname{Im}\theta_{\mu}$, where $\theta_{\mu}$ is the map in Theorem 2.6. Hence $\operatorname{Coker}(\sum_{c=1}^{{\ell}-1}\overline{\Psi}_{\mu(c)})\simeq L_{\mu}$ if and only if $\operatorname{Im}(\sum_{c=1}^{{\ell}-1}\overline{\Psi}_{\mu(c)})=\operatorname{Im}\theta_{\mu}$.
Again by Remark 3.10,
$\operatorname{Im}({\Psi}_{\mu(c)})\subseteq\operatorname{Im}(\Lambda^{\mu_{c}+1}\otimes\Lambda^{\mu_{c+1}-1}\xrightarrow{\theta_{1}}\Lambda^{\mu_{c}}\otimes\Lambda^{\mu_{c+1}}).$ (3.2)
@@
4 Schur modules and symmetrization of Garnir relations
@@
It is straightforward to verify that if $x\otimes y=e_{i_{1}}\cdots e_{i_{a}}\otimes e_{j_{1}}\cdots e_{j_{b}}$ is a basis element of $\Lambda^{a}\otimes\Lambda^{b}$, then $\phi_{k}(x\otimes y)$ is obtained by exchanging any $k$ elements $x_{1}<x_{2}<\dots<x_{k}$ of $\{e_{i_{1}},\dots,e_{i_{a}}\}$ with any $k$ elements $y_{1}<y_{2}<\dots<y_{k}$ of $\{e_{j_{1}},\dots,e_{j_{b}}\}$ and putting the $x_{1},\dots,x_{k}$ (respectively, $y_{1},\dots,y_{k}$) in the original positions of the $y_{1},\dots,y_{k}$ (respectively, $x_{1},\dots,x_{k}$).
@@
$\Phi_{t}=\sum_{i=0}^{t}(-1)^{t-i}\tbinom{b-t+i}{i}\phi_{t-i}=\sum_{k=0}^{t}(-1)^{k}\tbinom{b-k}{b-t}\phi_{k},$ (4.1)
where in the last equality we set $k=t-i$. Consider the $(b+1)\times(b+1)$ matrix $A=(a_{kt})$, where $a_{kt}=(-1)^{k}\tbinom{b-k}{b-t}$ and $0\leq k,t\leq b$. From the binomial identity $\sum_{j=0}^{b}(-1)^{k+j}\tbinom{b-k}{b-j}\tbinom{b-j}{b-t}=\delta_{k,t}$, see [9, p. 4], where $\delta_{k,t}$ is Kronecker’s delta, it follows that $A^{2}=I$, the identity matrix. Hence we may invert (4.1) to obtain $\phi_{k}=\sum_{t=0}^{k}(-1)^{t}\tbinom{b-t}{b-k}\Phi_{t}$ as desired.
The second equality follows from the first and Definition 4.13. ∎
@@
for all $j=1,\dots,b$.
5 Schur modules and Garnir relations